OpenStudy (akohl):

A curved section of roadway is in the shape of a circular arc of 410-m radius. The curved section is horizontal, not banked. The coefficient of friction for tires and the roadway is 0.40. What is the maximum safe driving speed for this unbanked, curved section of roadway?

OpenStudy (abhisar):

U can tag me in ur question like this @abhisar . It becomes easy for me to find uh :)

OpenStudy (abhisar):

It also sends me a notification !

OpenStudy (abhisar):

In this case the centripetal force is provided by only the frictional force.

OpenStudy (abhisar):

Now can u tell me the formula for centripetal force ?

OpenStudy (akohl):

Mass times velocity squared divided by the radius?

OpenStudy (abhisar):

good !

OpenStudy (akohl):

I'll start tagging you in my questions

OpenStudy (abhisar):

Now equate it with frictional force and find the value of velocity

OpenStudy (akohl):

What's the frictional force. All we have is the radius

OpenStudy (abhisar):

\(\mu \) = coefficient of friction.

OpenStudy (abhisar):

\(\color{blue}{\text{Originally Posted by}}\) @Abhisar \(\huge\sf \frac{\cancel mv^2}{r}=\mu \cancel mg\) \(\color{blue}{\text{End of Quote}}\)

OpenStudy (abhisar):

\(\color{blue}{\text{Originally Posted by}}\) @Abhisar ----> \(\huge\sf \frac{v^2}{r}=\mu g\) \(\color{blue}{\text{End of Quote}}\)

OpenStudy (akohl):

Thanks